OpenStudy (anonymous):

Which rates are equal? Choose 2. A. $57 per 20 boxes of rolls B. $45 per 18 rolls of paper towels C. $42 per 12 steaks D. $28 per 8 salads E. $26 per 10 bottles of soda

OpenStudy (kira_yamato):

A: 57/20 = 2.85 B: 45/18 = 2.5 C: 42/12 = 3.5 D: 28/8 = 3.5 E: 26/10 = 2.6 Therefore, C = D

OpenStudy (anonymous):

You'd have to divide each rate(: So 57 divided by 20, get the answer, and compare it to the others until you find the same rate(: I'll help you with the first one. 57 / 20 = 2.85 Now do that with the rest to see which ones have the same answer.

OpenStudy (anonymous):

thanks. @Kira_Yamato & @Yuba

OpenStudy (anonymous):

No problem, lol good luck.